# Viability Constraints on Baryogenesis in $f(R,L_{m},T)$ Gravity

### v1

@@ V Conclusion
A central unresolved conundrum in cosmology is the observed predominance of matter in the cosmos [8]. Among the proposed mechanisms to resolve this asymmetry is gravitational baryogenesis, which provides a possible explanation for the baryon–antibaryon imbalance [16]. This study investigates the phenomenon within the $f(R,L_{m},T)$ framework. Earlier studies provide strong motivation for this kind of exploration. For instance, the $f(R,L_{m},T)$ model satisfies energy criteria and admits stability for cosmological expansion of power-law form [30]. Furthermore, author in [31] suggested that $f(R,L_{m},T)$ model provides a viable setting to address broader challenges in cosmology.
The creation of matter imbalance in the cosmos can be attributed to charge-parity ($\mathcal{CP}$)-violating interactions, involving terms like $\partial_{\mu}(R+L_{m}+T)$ and $\partial_{\mu}f(R,L_{m},T)$, which acts as key operators for this mechanism. We explored four distinct cosmological eras under the assumption of a power-law form cosmic evolution. Our analysis produced BnER ratios $\frac{\eta_{B}}{s}$ in agreement with observationally established bounds, $\frac{\eta}{s}\cong 9.42\times 10^{-11}$. This outcome demonstrates that the $f(R,L_{m},T)$ model provides an effective paradigm for the observed asymmetry. In addition, generalized gravitational baryogenesis is also examined by modifying the interaction term to introduce extra degrees of freedom. The results show compatibility with observational constraints.
@@
    The expression for BnER within the framework of $f(R,L_{m},T)$ model is given by (27). In this work, we have examined four key epochs of cosmic evolution. By restricting the asymmetry ratio $\frac{\eta}{s}$ to the observationally acceptable interval, $7\times 10^{-11}\leq\eta/s\leq 2\times 10^{-10}$, we derive viable parameter spaces. By setting $\gamma=1$ and confining $\beta$ to the interval $10^{3}\;\leq\;\beta\;\leq\;-10^{2},$ the viable ranges of $\alpha$ are found to be:
    • Matter-dominated era ($\omega=0$):     $1.25\times 10^{156}\leq\alpha\leq 2.0\times 10^{159}.$
    • Radiation-dominated era ($\omega=1/3$): $2.25\times 10^{155}\leq\alpha\leq 3.0\times 10^{158}.$
    • Stiff-fluid era ($\omega=1$):         $7.75\times 10^{155}\leq\alpha\leq 9.0\times 10^{158}.$
    Furthermore, for the Quintessence era, characterized by $0.95\;\leq\;\omega\;\leq\;-0.35,$ we find two viable scenarios:
    1. 1.
      With $\gamma=1$ and $\beta=-5\times 10^{2}$, the parameter $\alpha$ is constrained to: $6.25\times 10^{153}\leq\alpha\leq 1.5\times 10^{158}$.
    2. 2.
      With $\gamma=1$ and $\alpha=1\times 10^{156}$, bound for the parameter $\beta$ is: $7.5\times 10^{3}\leq\beta\leq-52.5$.
  • GGB for $f(R,L_{m},T)$ model:
    The expression for BnER within $f(R,L_{m},T)$ gravity model under the generalized case is given by Eq. (33). To ensure the predicted asymmetry parameter $\eta/s$ resides within the observationally acceptable range of $7\times 10^{-11}\leq\eta/s\leq 2\times 10^{-10}$, the model parameters must satisfy specific constraints across different cosmological eras. For the matter-dominated era ($\omega=0$), the parameters $\beta$ and $\gamma$ must adhere to the relation $\beta=-\gamma$. In this scenario, the generation of a viable asymmetry requires $\beta>50.24$ and $\gamma<-50.24$. However, there is no explicit upper bound is imposed on $\beta$ nor a lower bound on $\gamma$. By further restricting $\beta$ and $\gamma$ to the intervals $50.24<\beta<1000$ and $1000<\gamma<-50.24$, respectively, the viable range for the parameter $\alpha$ is found to be $4.25\times 10^{-30}\leq\alpha\leq 2.758\times 10^{-22}$. For the stiff-fluid epoch ($\omega=1$), the necessary parameter relationship is $\beta=2\gamma$, constrained by the condition: $\beta<-25.12$, $\gamma<-12.56.$. By restricting $\beta$ and $\gamma$ to the ranges $1000<\beta<-25.12$ and $1000<\gamma<-12.56$, the corresponding viable values for $\alpha$ are constrained to $1.5\times 10^{-32}\leq\alpha\leq 2\times 10^{-25}$.
@@
    For radiation-dominated case, fixing $\gamma=1$ and confining $\beta$ to the interval $1\times 10^{3}\leq\beta\leq-1\times 10^{2}$, the variable $\alpha$ should reside in the range $6.0\times 10^{163}\leq\alpha\leq 2.0\times 10^{170}$ to produce the observed baryon asymmetry.
    The mathematical complexity of the generalized BnER expression precludes a full analysis across a continuous range of the EoS parameter $\omega$. Therefore, for the quintessence era, we select a representative value of $\omega=-0.65$ for our analysis. For the generation of the asymmetry in this regime, the model parameters $\beta$ and $\gamma$ must satisfy the condition $\gamma=-0.34\beta$. Furthermore, the constraints require $\beta<-157,\gamma>54$. By further restricting $\beta$ and $\gamma$ to the intervals $1000<\beta<-157$ and $54<\gamma<1000$, respectively, the viable range for the parameter $\alpha$ is found to be $7.25\times 10^{-30}\leq\alpha\leq 8.5\times 10^{-23}$.
@@
Based on the obtained results, it can be implied that $f(R,L_{m},T)$ model offers a viable setting for realizing gravitational baryogenesis, under favoured model parameter bounds. It is an important point to note that the linear $f(R,L_{m},T)$ model is primarily an analytically tractable initial step to probe matter–geometry couplings in the baryogenesis context. Although the linear term in $R$ does not introduce any extra dynamical degrees of freedom beyond GR, the model’s novelty is mainly encoded in the explicit couplings to $L_{m}$ and $T$. This work can be extended to study different models of $f(R,L_{m},T)$ gravity with nonlinear extensions of $R$ under some other viable forms of scale factor.

### v2

@@ V Conclusion
A central unresolved conundrum in cosmology is the detected excess of matter in the cosmos [8]. Among the proposed mechanisms to resolve this asymmetry is gravitational baryogenesis, which provides a possible explanation for the baryon–antibaryon imbalance [16, 46]. This study investigates the phenomenon within the $f(R,L_{m},T)$ paradigm. Earlier studies provide strong motivation for this kind of exploration. For instance, the $f(R,L_{m},T)$ model satisfies energy criteria and admits stability for cosmological expansion of power-law form [31]. Furthermore, author in [32] suggested that $f(R,L_{m},T)$ model presents a viable setting to address broader challenges in cosmology.
The creation of matter imbalance in the cosmos can be attributed to charge-parity ($\mathcal{CP}$)-violating interactions, involving terms like $\partial_{\mu}(R+L_{m}+T)$ and $\partial_{\mu}f(R,L_{m},T)$, which acts as key operators for this mechanism. Our analysis produced BnER ratios $\frac{\eta_{B}}{s}$ in agreement with observationally established bounds, $\frac{\eta}{s}\cong 9.42\times 10^{-11}$. This outcome demonstrates that the $f(R,L_{m},T)$ model provides an effective explanation for the detected imbalance. Furthermore, generalized gravitational baryogenesis is also examined by modifying the interaction term to introduce extra degrees of freedom. The results show compatibility with observational constraints.
@@ V.1 Model I
    The expression for BnER for model I is given by (27). In this work, by restricting the asymmetry ratio $\frac{\eta}{s}$ to the observationally acceptable interval: $2\times 10^{-10}\geq\eta/s\geq 7\times 10^{-11}$, we derive a viable parameter space. By setting $\gamma=1$ and confining $\beta$ to the interval $10^{3}\;\leq\;\beta\;\leq\;-10^{2},$ the viable range of $\alpha$ is found to be $3.5\times 10^{38}\leq\alpha\leq 4.25\times 10^{41}.$
  • GGB in Model I:
@@
    The expression for BnER within $f(R,L_{m},T)$ gravity model I under the generalized case is given by Eq. (36). To ensure that the predicted asymmetry parameter $\eta/s$ lies within the observationally acceptable interval $2\times 10^{-10}\geq\eta/s\geq 7\times 10^{-11}$, by fixing $\gamma=1$ and confining $\beta$ to the interval $1\times 10^{3}\leq\beta\leq-1\times 10^{2}$, the variable $\alpha$ should reside in the range $3.5\times 10^{46}\leq\alpha\leq 4.5\times 10^{53}$ to produce the observed baryon asymmetry.
V.2 Model II
  • GB in Model II:
    The analytical form of BnER corresponding to Model II is presented by Eq. (32). In this analysis, the asymmetry ratio $\eta/s$ is constrained within the observationally supported range $2\times 10^{-10}\geq\eta/s\geq 7\times 10^{-11}$, ensuring compatibility with cosmological bounds. Under the choice $\gamma=1$ and restricting the parameter $\beta$ to the interval $1\times 10^{-32}\leq\beta\leq 1\times 10^{-27}$, the corresponding viable range of $\alpha$ is obtained as $5\times 10^{44}\leq\alpha\leq 2\times 10^{48}$.
  • GGB in Model II:
    The mathematical form of BnER for the gravity Model II under the generalized scenario is described by Eq. (38). In order to maintain consistency with the observationally permitted range of the asymmetry parameter, $7\times 10^{-11}\leq\eta/s\leq 2\times 10^{-10}$, the parameters are appropriately constrained. By setting $\gamma=1$ and restricting $\beta$ within $1\times 10^{-36}\leq\beta\leq 1\times 10^{-34}$, the parameter $\alpha$ is found to lie in the range $1\times 10^{46}\leq\alpha\leq 4.25\times 10^{56}$, which successfully yields the observed matter asymmetry.
Comparative analysis of the models reveals a systematic distinction in their coupling behavior. Both models operate in a regime of stronger curvature coupling ($\alpha$) of order $10^{38}-10^{56}$, showing significant departures from GR, where $\alpha=\frac{1}{2}$. Linear Model I requires a strong and negative matter coupling ($\beta\sim-10^{2}\text{ to }-10^{3}$), which presents a more natural parameter setting. In contrast, non-linear Model II requires an extremely weak but positive coupling ($\beta\sim 10^{-27}\text{ to }10^{-36}$). As a consequence, Model I emerges as the more theoretically plausible framework for generating the observed baryon asymmetry.
@@ V.2 Model II
Based on the obtained results, it can be implied that $f(R,L_{m},T)$ model offers a viable setting for realizing gravitational baryogenesis, under favoured model parameter bounds. This work can be extended to study different models of $f(R,L_{m},T)$ theory under some other feasible forms of scale factor.